2x ^3(−16x^3y^7) + 5y ^3(250x^6y^4)

Answer by Boreal(15235)

2x ^3(−16x^3y^7) + 5y ^3(250x^6y^4)=
-32x^6*y^7+1250(x^6)*(y^7)=1228*x^6*y^7